OpenStudy (anonymous):

-2sin^2(x)-1cos(x)=-1 Looking for the smallest non-negative solution and the next smallest non-negative solution Smallest non-negative solution: 0, ? Next smallest non-negative solution: 360, ?

OpenStudy (anonymous):

-2sin^2(x)-cos(x)=-1 -2(1-cos^2(x))-cos(x)=-1 -2+2cos^2(x)-cos(x)=-1 2cos^2(x)-cos(x)-1=0 (cosx-1)(2cosx+1)=0 cosx=1 or cosx=-1/2 x=0,360,720,... or x=120, 240, ... therefore, Smallest non-negative solution=0 and the next smallest non-negative solution =120